5.) Soar Incorporated is considering eliminating its mountain bike division, which reported an operating loss for the recent year of $6,000. The division sales for the year were $1,044,000 and the variable costs were $863,000. The fixed costs of the division were $187,000. If the mountain bike division is dropped, 30% of the fixed costs allocated to that division could be eliminated. The impact on operating income for eliminating this business segment would be: Multiple Choice $56,100 decrease $124,900 decrease $50,100 decrease $181,000 increase $181,000 decrease 6.) Walters manufactures a specialty food product that can currently be sold for $22.40 per unit and has 20,400 units on hand. Alternatively, it can be further processed at a cost of $12,400 and converted into 12,400 units of Deluxe and 6,400 units of Super. The selling price of Deluxe and Super are $30.40 and $20.40, respectively. The incremental net income of processing further would be: Multiple Choice $38,160. $50,560. $18,400. $44,400. $12,400. 7.) After-tax net income divided by the average amount invested in a project, is the: Multiple Choice Net present value rate. Payback rate. Accounting rate of return. Earnings from investment. Profit rate. 8.) A new manufacturing machine is expected to cost $278,000, have an eight-year life, and a $30,000 salvage value. The machine will yield an annual incremental after-tax income of $35,000 after deducting the straight-line depreciation. Compute the payback period for the purchase. Multiple Choice 8.7 years. 3.8 years. 4.2 years. 7.3 years. 5.4 years. 9.) A company is considering purchasing a machine for $21,000. The machine will generate an after-tax net income of $2,000 per year. Annual depreciation expense would be $1,500. What is the payback period for the new machine? Multiple Choice 4 years. 6 years. 10.5 years. 14 years. 42 years.
Answer #1

Impact on operating income = Fixed costs avoided – contribution margin lost

= 187,000*30% - (1,044,000-863,000)

= $124,900 decrease

6.iNcremental income from processing = Sales after processing – sales if not processed – cost of processing

= 12,400*30.40 + 6,400*20.40 – 20,400*22.4 – 12,400

= $38,160

7.Accounting rate of return = Net income/Average Investment

8.Payback period = Initial cost/Annual cash flows

= 278,000/(35,000+31,000)

= 4.21 years

Annual depreciation = (Cost – salvage value)/life

9.Payback period = 21,000/(2,000+1,500)

= 6 years